In a photoelectric experiment, the collector plate is at 2.0 V with respect to the emitter plate made of copper (φ = 4.5 eV). The emitter is illuminated by a source of monochromatic light of wavelength 200 nm. Find the minimum and maximum kinetic energy of the photoelectrons reaching the collector.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. ∅ = 4.5 eV, λ = 200 nm Stopping potential or energy = E - ∅ = WC/λ - ∅ Minimum 1.7 V is necessary to stop the electron The minimum K.E. = 2eV [Since the electric potential of 2 V is reqd. to accelerate the electron to reach the plates] the maximum K.E. = (2+1, 7)ev = 3.7 ev.